# OpeNLGauge: An Explainable Metric for NLG Evaluation with Open-Weights LLMs

### v1

@@ A.5.1 Datasets
  • NarrativeQA (Kočiský et al., 2018) consists of human-written questions and free-form answers based on stories or their summaries. The stories include books and movie scripts and are provided either as full texts or as human-written summaries. Since full stories do not fit into the context window of many LLMs, we use only summaries to generate the outputs. Along with newly generated answers, we include human reference answers from the original datasets in the evaluated outputs.
  • FeTaQA (Nan et al., 2022b) is a question answering dataset based on Wikipedia tables that requires models to aggregate and reason about the entities in the table and their relations. This places the task at the intersection of data-to-text and question answering task categories. In addition to new LLM outputs, we use pre-generated outputs from Zhao et al. (2023b).
@@ A.5.2 Aspects
  • Correctness evaluates if the answer to a question is correct with respect to the input. Since our models are instructed to assess the quality on an ordinal scale, we evaluate a degree of correctness - the answer should receive the maximum score if it is fully correct, while lower scores should reflect the number and severity of correctness issues. We define correctness as: Extent to which the answer to the question is correct with respect to the input.
@@
  • Relevance is concerned with specificity of the answer and measures the degree to which the answer addresses the particular question asked. Although it is related to conciseness, relevance is not that much concerned with the amount of detail in the answer. We define relevance as: Extent to which the answer is specific and meaningful with respect to the question.
@@ Appendix B Collected system outputs
Table 6: Overview of the evaluated systems in the dataset for the summarization task. Sources: GW = Gao and Wan (2022), GR = Grusky et al. (2018), NA = Narayan et al. (2018), ST = Stiennon et al. (2020), new = newly generated.
@@
Qwen 2.5 0.5B (Yang et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
@@
T5 (Raffel et al., 2020) pre-trained LM ST
UniLM (Dong et al., 2019) pre-trained LM ST
CODS (Wu et al., 2021) BART-based hybrid model GW
ConvoSumm (Fabbri et al., 2021a) BART-based model GW
Ctrl-DiaSumm (Liu and Chen, 2021) BART-based model GW
ConvS2S (Gehring et al., 2017) Convolutional seq-to-seq NA
Topic-ConvS2S (Narayan et al., 2018) Topic-conditioned ConvS2S NA
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq with attention GR
PNG (Vinyals et al., 2015; Gulçehre et al., 2016) Pointer-generator network GR, GW, NA
TextRank (Barrios et al., 2016) Ranking-based extractive summarization GR
@@
LEAD-3 (See et al., 2017) Extractive baseline GR, GW, ST
@@
Table 7: Overview of the evaluated systems in the dataset for the data-to-text task. Sources: DU = Dušek et al. (2020), FE = Ferreira et al. (2020), ZH = Zhao et al. (2023b), new = newly generated.
@@
Qwen 2.5 Coder 1.5B (Yang et al., 2024) Instruction-tuned LLM new
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Solar 10.7B (Kim et al., 2024a) Instruction-tuned LLM new
DeepSeek Coder v2 16B (Zhu et al., 2024) Instruction-tuned LLM new
@@
NILC (Sobrevilla Cabezudo and Pardo, 2020) Fine-tuned BART FE
Orange-NLG (Montella et al., 2020) Fine-tuned BART FE
Amazaon AI (Shanghai) (Guo et al., 2020) Graph CNN + T5 FE
GPT2-C2F (Chen et al., 2020) Fine-tuned GPT-2 ZH
LoFT (Zhao et al., 2023a) Fine-tuned BART ZH
PLOG (Liu et al., 2022a) Fine-tuned T5 ZH
R2D2 (Nan et al., 2022a) Fine-tuned T5 ZH
Flan-T5 (Chung et al., 2024) Instruction-tuned LM ZH
Adapt (Elder et al., 2018) RNN seq-to-seq DU
Sheff2 (Chen et al., 2018) RNN seq-to-seq DU
Slug (Juraska et al., 2018) RNN + convolutional seq-to-seq DU
Forge1 (Mille and Dasiopoulou, 2018) Rule-based DU
TR2 (Smiley et al., 2018) Template-based DU
DANGNT-SGU (Tran and Nguyen, 2020) Template-based FE
RALI-Université de Montréal (Lapalme, 2020) Template-based FE
Table 8: Overview of the evaluated systems in the dataset for the dialogue response generation task. Sources: GU = Gupta et al. (2019), HU = Huang et al. (2020), ZH = Zhao et al. (2020), new = newly generated.
@@
Tülu 3 (Lambert et al., 2024) Instruction-tuned LLM new
@@
Vicuna 7B (Zheng et al., 2023) Instruction-tuned LLM new
BlenderBot-small (Roller et al., 2021) Dialogue LM new
DialoGPT-small (Zhang et al., 2020) Dialogue LM new
GPT-Neo 125M (Gao et al., 2020) Pre-trained LM new
GPT-2 (Wolf et al., 2019) Pre-trained LM ZH
CVAE (Zhao et al., 2017) Conditional variational autoencoder GU
HRED (Serban et al., 2016) Hierarchical recurrent encoder-decoder GU, ZH
Transformer-generator (Dinan et al., 2018) Transformer-based generative model HU
Transformer-ranker (Urbanek et al., 2019) Transformer-based ranking model HU
DualEncoder (Lowe et al., 2015) LSTM dual encoder GU
VHRED (Serban et al., 2017) Latent variable HRED ZH
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq with attention GU, ZH
Table 9: Overview of the evaluated systems in the dataset for the story generation task. Sources: GU = Guan et al. (2021), new = newly generated.
@@
Gemma 2 2B (Gemma Team et al., 2024) Instruction-tuned LLM new
@@
GPT-2 (Radford et al., 2019) Pre-trained LM GU
GPT-KG (Guan et al., 2020) Knowledge-enhanced GPT-2 GU
Fusion (Fan et al., 2018) Convolutional seq-to-seq with attention GU
Plan&Write (Yao et al., 2019) Hierarchical RNN-based model GU
S2S (Cho et al., 2014; Sutskever et al., 2014) RNN-based seq-to-seq GU
Table 10: Overview of the evaluated systems in the dataset for the question answering task. Sources: KO = Kočiský et al. (2018), ZH = Zhao et al. (2023b), new = newly generated.
@@
DeepSeek Coder v2 16B (Zhu et al., 2024) Instruction-tuned LLM new
Llama 2 7B Chat (Touvron et al., 2023) Instruction-tuned LLM new
Qwen 2.5 Coder 1.5B (Yang et al., 2024) Instruction-tuned LLM new
Qwen 2.5 0.5B (Yang et al., 2024) Instruction-tuned LLM new
GPT-Neo 125M (Gao et al., 2020) Pre-trained LM new
BART (Lewis, 2019) Pre-trained LM ZH
Flan-T5 (Chung et al., 2024) Instruction-tuned LM ZH
OmniTab (Jiang et al., 2022) Table pre-trained LM ZH
ReasTAP (Zhao et al., 2022) Table pre-trained LM ZH
TAPEX (Liu et al., 2022b) Table pre-trained LM ZH
@@ Appendix C LLMs used in the experiments
Qwen 2.5 70B 8-bit qwen2.5:72b-instruct-q8_0
@@
Appendix D Meta-evaluation datasets
@@
We used the following datasets for meta-evaluation:
@@
  • SummEval (Fabbri et al., 2021b) is a standard meta-evaluation dataset for summarization. It consists of summaries generated from CNN/DailyMail articles, with 100 input articles and 16 different system outputs for each article. Human evaluations address four aspects: consistency, relevance, coherence, and fluency. Each output is scored by three expert annotators and five crowdworkers. We use only averaged expert annotations.
@@
  • QAGS (Wang et al., 2020a) contains annotations of consistency for summaries from CNN/DailyMail and XSum datasets. It includes 235 CNN/DailyMail summaries and 239 XSum summaries, each annotated by three evaluators. Annotators assign a binary factual consistency score (yes/no) for each sentence of the summary. We follow (Wang et al., 2020a) and apply majority voting for each sentence annotation, followed by averaging sentence-level scores to obtain the overall score for the summary.
@@
  • HANNA (Chhun et al., 2022) is a benchmark for story generation based on the WritingPrompts dataset. It contains human evaluations for 1056 stories across six aspects: relevance, coherence, engagement, empathy, surprise and complexity (see Appendix A.4 for details).
@@
  • SFRES and SFHOT (Wen et al., 2015) are used to perform a meta-evaluation for data-to-text. These datasets consist of dialogue acts (DAs) in structured format with generated responses providing information about restaurants and hotels in San Francisco. Human judgments are provided for informativeness and naturalness.
@@
  • TopicalChat (Gopalakrishnan et al., 2023) annotations from the USR dataset (Mehri and Eskenazi, 2020b) are used for meta-evaluation on dialogue tasks. The dataset includes human evaluations for five aspects: groundedness, coherence, interestingness, naturalness and understandability. As the source data differ from our training data, TopicalChat serves as out-of-domain evaluation dataset. Additionally, since groundedness is not present in our training data, we evaluate it as an unseen aspect.
@@
  • Wiki-DA (Alva-Manchego et al., 2021) is a dataset of DA human ratings for text simplification, an unseen task for OpeNLGaugeft during training. Along with scores for fluency, this dataset also includes two unseen aspects: meaning preservation and simplicity.
Appendix E Prompt Templates
@@
Prompt templates for annotator and supervisor LLMs are presented in Listings 12. Prompt template used for fine-tuning and inference of the distilled model is shown in Listing 3.
Appendix F Full Results
@@
Tables 1218 contain the meta-evaluation results for additional datasets described in Appendix D.
Appendix G Ablation experiments
G.1 Prompt ablations
@@
G.2 Ensemble size ablations
The effect of the ensemble size on Spearman’s correlations are presented in Figures 49.
Appendix H Output Examples

### v3

@@ A.5.1 Datasets
  • NarrativeQA (kovcisky2018narrativeqa) consists of human-written questions and free-form answers based on stories or their summaries. The stories include books and movie scripts and are provided either as full texts or as human-written summaries. Since full stories do not fit into the context window of many LLMs, we use only summaries to generate the outputs. Along with newly generated answers, we include human reference answers from the original datasets in the evaluated outputs.
  • FeTaQA (nan2022fetaqa) is a question answering dataset based on Wikipedia tables that requires models to aggregate and reason about the entities in the table and their relations. This places the task at the intersection of data-to-text and question answering task categories. In addition to new LLM outputs, we use pre-generated outputs from zhao-etal-2023-investigating.
@@ A.5.2 Aspects
  • Correctness evaluates if the answer to a question is correct with respect to the input. Since our models are instructed to assess the quality on an ordinal scale, we evaluate a degree of correctness – the answer should receive the maximum score if it is fully correct, while lower scores should reflect the number and severity of correctness issues. We define correctness as: Extent to which the answer to the question is correct with respect to the input.
@@
  • Relevance is concerned with the specificity of an answer and measures the degree to which the answer addresses the particular question asked. Although it is related to conciseness, relevance is not that much concerned with the amount of detail in the answer. We define relevance as: Extent to which the answer is specific and meaningful with respect to the question.
@@ Appendix B Collected system outputs
Table 6: Overview of the evaluated systems in the dataset for the summarization task. Sources: GW = gao2022dialsummeval, GR = grusky2018newsroom, NA = narayan2018don, ST = stiennon2020learning, new = newly generated.
@@
Qwen 2.5 0.5B (yang2024qwen2) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
@@
T5 (raffel2020exploring) pre-trained LM ST
UniLM (dong2019unified) pre-trained LM ST
CODS (wu-etal-2021-controllable) BART-based hybrid model GW
ConvoSumm (fabbri-etal-2021-convosumm) BART-based model GW
Ctrl-DiaSumm (liu-chen-2021-controllable) BART-based model GW
ConvS2S (gehring2017convolutional) Convolutional seq-to-seq NA
Topic-ConvS2S (narayan2018don) Topic-conditioned ConvS2S NA
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq with attention GR
PNG (vinyals2015pointer; gu̇lccehre2016pointing) Pointer-generator network GR, GW, NA
TextRank (barrios2016variations) Ranking-based extractive summarization GR
@@
LEAD-3 (see2017get) Extractive baseline GR, GW, ST
@@
Table 7: Overview of the evaluated systems in the dataset for the data-to-text task. Sources: DU = duvsek2020evaluating, FE = ferreira20202020, ZH = zhao-etal-2023-investigating, new = newly generated.
@@
Qwen 2.5 Coder 1.5B (yang2024qwen2) Instruction-tuned LLM new
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Solar 10.7B (kim-etal-2024-solar) Instruction-tuned LLM new
DeepSeek Coder v2 16B (zhu2024deepseek) Instruction-tuned LLM new
@@
NILC (sobrevilla-cabezudo-pardo-2020-nilc-webnlg) Fine-tuned BART FE
Orange-NLG (montella2020denoising) Fine-tuned BART FE
Amazaon AI (Shanghai) (guo-etal-2020-2) Graph CNN + T5 FE
GPT2-C2F (chen2020logical) Fine-tuned GPT-2 ZH
LoFT (zhao2023loft) Fine-tuned BART ZH
PLOG (liu-etal-2022-plog) Fine-tuned T5 ZH
R2D2 (nan2022r2d2) Fine-tuned T5 ZH
Flan-T5 (chung2024scaling) Instruction-tuned LM ZH
Adapt (elder-etal-2018-e2e) RNN seq-to-seq DU
Sheff2 (chen2018sheffield) RNN seq-to-seq DU
Slug (juraska-etal-2018-deep) RNN + convolutional seq-to-seq DU
Forge1 (mille2018forge) Rule-based DU
TR2 (smiley-etal-2018-e2e) Template-based DU
DANGNT-SGU (tran-nguyen-2020-webnlg) Template-based FE
RALI-Université de Montréal (lapalme-2020-rdfjsrealb) Template-based FE
Table 8: Overview of the evaluated systems in the dataset for the dialogue response generation task. Sources: GU = gupta2019investigating, HU = huang2020grade, ZH = zhao2020designing, new = newly generated.
@@
Tülu 3 (lambert2024tulu) Instruction-tuned LLM new
@@
Vicuna 7B (zheng2023judging) Instruction-tuned LLM new
BlenderBot-small (roller-etal-2021-recipes) Dialogue LM new
DialoGPT-small (zhang-etal-2020-dialogpt) Dialogue LM new
GPT-Neo 125M (gao2020pile) Pre-trained LM new
GPT-2 (wolf2019transfertransfo) Pre-trained LM ZH
CVAE (zhao2017learning) Conditional variational autoencoder GU
HRED (serban2016building) Hierarchical recurrent encoder-decoder GU, ZH
Transformer-generator (dinan2018wizard) Transformer-based generative model HU
Transformer-ranker (urbanek-etal-2019-learning) Transformer-based ranking model HU
DualEncoder (lowe2015ubuntu) LSTM dual encoder GU
VHRED (serban2017hierarchical) Latent variable HRED ZH
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq with attention GU, ZH
Table 9: Overview of the evaluated systems in the dataset for the story generation task. Sources: GU = guan2021openmeva, new = newly generated.
@@
Gemma 2 2B (gemma2024gemma) Instruction-tuned LLM new
@@
GPT-2 (radford2019language) Pre-trained LM GU
GPT-KG (guan2020knowledge) Knowledge-enhanced GPT-2 GU
Fusion (fan2018hierarchical) Convolutional seq-to-seq with attention GU
Plan&Write (yao2019plan) Hierarchical RNN-based model GU
S2S (cho2014learning; sutskever2014sequence) RNN-based seq-to-seq GU
Table 10: Overview of the evaluated systems in the dataset for the question answering task. Sources: KO = kovcisky2018narrativeqa, ZH = zhao-etal-2023-investigating, new = newly generated.
@@
DeepSeek Coder v2 16B (zhu2024deepseek) Instruction-tuned LLM new
Llama 2 7B Chat (touvron2023llama) Instruction-tuned LLM new
Qwen 2.5 Coder 1.5B (yang2024qwen2) Instruction-tuned LLM new
Qwen 2.5 0.5B (yang2024qwen2) Instruction-tuned LLM new
GPT-Neo 125M (gao2020pile) Pre-trained LM new
BART (lewis2019bart) Pre-trained LM ZH
Flan-T5 (chung2024scaling) Instruction-tuned LM ZH
OmniTab (jiang-etal-2022-omnitab) Table pre-trained LM ZH
ReasTAP (zhao-etal-2022-reastap) Table pre-trained LM ZH
TAPEX (liu2022tapex) Table pre-trained LM ZH
@@ Appendix C LLMs used in the experiments
Qwen 2.5 72B 8-bit qwen2.5:72b-instruct-q8_0
@@
Appendix D Outliers
Table 12 shows the percentages of outliers detected for each annotator model in the ensemble and task category during synthetic data generation. To keep the merged evaluation outputs internally consistent, these outliers were removed before merging the outputs of the individual LLMs.
The proportions in the table indicate that the final annotation utilized most of the generated evaluations from the individual LLMs. The exceptions include Command R+ 104B and Gemma 2 27B on data-to-text tasks, with 29% and 10% outliers, respectively. Additionally, 10% evaluation outputs of Command R+ 104B were removed for question answering tasks. However, on average only 3.4% of evaluation outputs are detected as outliers across all model-task pairs.
Model Data-to-text Summarization Dialogue Story Generation Question Answering
Command R+ 104B 28.88 0.13 0.73 0.94 9.61
Gemma 2 27B 10.32 1.96 1.52 1.03 5.07
Llama 3.1 Nemotron 70B 4.27 1.48 4.02 1.03 0.85
Mistral Large 2 123B 4.89 0.93 0.63 0.25 1.22
Qwen 2.5 72B 1.39 1.36 1.43 1.12 0.47
Table 12: Proportions (%) of evaluation outputs detected as outliers for each model and task category. Across all model-task pairs, 3.4% of evaluation outputs are detected as outliers on average (median = 1.36%).
Appendix E Meta-evaluation datasets
@@
We used the following datasets for meta-evaluation. In the datasets where human scores are not already aggregated, we averaged the scores of all annotators.
@@
  • SummEval (fabbri2021summeval) is a standard meta-evaluation dataset for summarization. It consists of summaries generated from CNN/DailyMail articles, with 100 input articles and 16 different system outputs for each article. Human evaluations address four aspects: (factual) consistency, relevance, coherence, and fluency. Each output is scored by three expert annotators and five crowdworkers. We use only expert scores to measure correlations.
@@
  • QAGS (wang2020asking) contains annotations of consistency for summaries from CNN/DailyMail and XSum datasets. It includes 235 CNN/DailyMail summaries and 239 XSum summaries, each annotated by three evaluators. Annotators assign a binary factual consistency score (yes/no) for each sentence of the summary. We follow (wang2020asking) and apply majority voting for each sentence annotation, followed by averaging sentence-level scores to obtain the overall score for the summary.
@@
  • HANNA (chhun2022human) is a benchmark for story generation based on the WritingPrompts dataset. It contains three human scores for each of the 1056 stories across six aspects: relevance, coherence, engagement, empathy, surprise and complexity (see Appendix A.4 for details on these aspects).
@@
  • SFRES and SFHOT (wen2015semantically) are used to perform a meta-evaluation for data-to-text. These datasets consist of dialogue acts (DAs) in structured format with generated responses providing information about restaurants and hotels in San Francisco. Human judgments are provided for informativeness and naturalness.
@@
  • TopicalChat (gopalakrishnan2023topical) annotations from the USR dataset (mehri2020usr) are used for meta-evaluation of dialogue response generation. The dataset includes human evaluations for five aspects: groundedness, coherence, interestingness, naturalness and understandability. As the source data differ from our training data, TopicalChat serves as an out-of-domain evaluation dataset. Additionally, since groundedness is not present in our training data, we evaluate it as an unseen aspect.
@@
  • Wiki-DA (alva2021suitability) is a dataset of DA human ratings for text simplification, a task unseen by OpeNLGaugeft during training. Along with scores for fluency, this dataset also includes two unseen aspects: meaning preservation and simplicity.
Appendix F Prompt templates
@@
Prompt templates for annotator and consolidator LLMs are presented in Listings 12. Prompt template used for fine-tuning and inference of the distilled model is shown in Listing 3.
Appendix G Full results
System
Error span
Error
No span given
Not an error
Hallucination
Total
Explanation
OpeNLGaugeens
Correct 218 9 3 0 230
Partially correct 41 0 0 1 42
Incomplete 20 0 1 0 21
Vague 4 2 1 0 7
Incorrect 20 3 32 1 56
Not an error 0 0 10 0 10
Total 303 14 47 2 366
Span OK (%) 83
Exp. correct (%) 63
OpeNLGaugeft
Correct 108 3 4 4 119
Partially correct 60 1 2 0 63
Incomplete 9 0 0 0 9
Vague 3 9 0 0 12
Incorrect 41 7 42 6 96
Not an error 0 1 8 0 9
Total 221 21 56 10 308
Span OK (%) 72
Exp. correct (%) 39
TigerScore
Correct 8 17 0 0 25
Partially correct 5 2 0 0 7
Incomplete 2 2 0 0 4
Vague 0 3 0 0 3
Incorrect 5 16 2 0 23
Not an error 1 2 0 0 3
Total 21 42 2 0 65
Span OK (%) 32
Exp. correct (%) 38
Table 13: Detailed results of human evaluation of error-spans and their explanations. Each table shows absolute occurrence counts of different error span and explanation validity (see Section 7.2), with overall proportions of correct annotation given below. The reported differences between TigerScore and OpeNLGaugeens are statistically significant (t-test, $p<0.05$).
@@ Appendix G Full results
Tables 1420 contain the meta-evaluation results for additional datasets described in Appendix E. Detailed results of human evaluation of error span quality are presented in Table 13.
G.1 Kendall’s $\tau$ correlations
On QAGS, our method achieves lower Kendall’s $\tau$ than Themis, although it surpasses the metric on Spearman’s $\rho$. This discrepancy can be attributed to a difference in score precisions between these two methods. As a result of averaging, OpeNLGaugeens provides more granular floating point scores, while Themis predicts integer scores on a Likert scale. Generally, we can expect more tied pairs (i.e. pairs that are neither concordant nor discordant) in the calculation of Kendall’s $\tau$ with integer scores, which can have a substantial effect on the correlations. In QAGS, 78% of human scores map to integer scores after normalization. Therefore, we consider Spearman’s $\rho$ more appropriate for comparing evaluation capabilities of different metrics.
Similarly, specific individual LLMs score higher on Kendall’s $\tau$ than the ensemble on some datasets (see Tables 3, 14, 18 and 19). This is due to the same reason outlined above. Specifically, there are 37% tied pairs for OpeNLGaugeens and human scores, while the individual LLMs, which provide integer scores, show substantially larger proportions of tied pairs: between 46% and 54%.
Appendix H Ablation experiments
H.1 Prompt ablations
@@
H.2 Ensemble size ablations
The effect of the ensemble size on Spearman correlations are presented in Figures 510.
Appendix I Inter-annotator agreement between ensemble models
To obtain additional insights into the variance between individual ensemble LLMs in their predicted overall scores, we compute pairwise inter-annotator agreements between the models on all meta-evaluation datasets.
Figure 16 shows the Spearman correlations for each LLM pair and dataset used in our meta-evaluation, where the correlations are calculated over all evaluation aspects in a given dataset. The agreement is generally high on the QAGS and Wiki-DA datasets, although substantially lower on other datasets. Note that individual models also achieved relatively high correlations with humans on these two datasets, which indicates that they might generally consist of examples (and possibly evaluation aspects) for which it is easier for LLMs to make decisions on their quality. In contrast, the models show relatively low agreement on HANNA, which could be explained by the more subjective nature of the story generation task and the corresponding evaluation aspects. [^21] While such hypotheses could in principle be tested by comparing our results with the agreement between human annotators on a particular task, there are several issues that limit the reliability of such comparisons – for each meta-evaluation dataset we use, at least two of the following hold: (1) the number of human annotators is different from the number of LLMs in our ensemble, (2) either the sets of human annotators differ between evaluated outputs, or it is not made explicit in the dataset (or the corresponding paper) whether this is the case, (3) rating scales are different from ours (often three levels, or even binary). Across all tasks, Command R+ 104B tends to disagree more with other LLMs when compared to other pairwise agreements.
To assess the agreement of the models on exact overall score predictions, we also measure pairwise Cohen’s $\kappa$ (Figure 17), which treats the overall scores as categorical and therefore serves as a stricter measure. As expected, the agreements are lower compared to those measured by Spearman correlations, while the general trend is the same: the agreement on exact scores is higher on QAGS and Wiki-DA than on other datasets, while Command R+ 104B generally disagrees more with other LLMs across tasks.
Finally, we compute Krippendorff’s $\alpha$ (krippendorff2011computing), which allows us to measure the agreement between all ensemble models, while also being applicable to ordinal data. In general, the agreements are low to moderate. Similarly to the pairwise agreements, we observe substantially higher agreements on QAGS and Wiki-DA, while the models agree the least on scoring generated stories in the HANNA dataset.
Overall, our analysis indicates that the predictions of individual models are sufficiently diverse to benefit from the combination in an ensemble without much redundancy.
Appendix J Score aggregation methods
Table 21 compares results of different approaches to aggregation of ensemble scores:
  • Average computes the final score for the evaluated output as a simple average of scores from all individual LLMs.
  • Average w/o outliers first removes the outliers before computing the average. The score is considered an outlier if it differs from the average of other scores for the same example by at least two standard deviations and this difference is at least 1.
  • Median computes the final score as a median of the scores from individual LLMs to disregard the effect of extreme scores.
  • Majority voting selects the most frequent score from the individual LLMs for the given output.
  • Min selects the minimum of individual scores as the final score. This corresponds to the most strict evaluation, i.e. typically the evaluation that detected the most errors or assigned highest severity levels to the identified errors.
Appendix K False positives analysis
To estimate the extent of potential over-annotation by OpeNLGaugeens and its components, we analyze overall score predictions for output-aspect pairs $(y,a)$ which obtained maximum scores by all human annotators in a given dataset. We denote these examples $Y_{max}$ and assume that if $(y,a)\in Y_{max}$, then $y$ does not contain any errors with respect to aspect $a$.
[…]
Figure 18 shows the distribution of error counts (top), mean severity levels (middle) and maximum severity levels (bottom) per output, as predicted by OpeNLGaugeens and its components for the $Y_{max}$ subset of SummEval. Most individual LLMs assess the outputs in $Y_{max}$ as error-free, particularly Command R+ 104B, which agrees almost perfectly with human annotators in this subset. Note that in general, Command R+ 104B tends to disagree the most with other LLMs, as discussed in Appendix I, while also achieving the highest correlations with humans in evaluating factual consistency (Table 14), which represents approximately half of the examples in $Y_{max}$. In contrast, the ensemble shows a tendency to include at least one error in most of its annotations for $Y_{max}$. This could be attributed to the merging procedure by the consolidator model, which aggregates errors from five different models and could lead to accumulation of errors. As expected, an increasing number of errors detected by a model is also reflected in the mean and maximum severity levels.
The error annotations for TopicalChat (Figure 19) show a similar pattern, although a larger proportion of individual models have a tendency to detect one or more errors in this case.
Appendix L Output examples